1. Given the following supply and demand model:
QD = a − bP + eY0 Demand
QS = −c + dP + fPr Supply
QD = QS Equilibrium
where: a, b, c, d, e > 0, Y0 is exogenous income, Pr is the exogenous price of a related good.
a) What are the necessary conditions for positive equilibrium prices and quantities?
b) What is the economic interpretation of the parameter "f"?
c) What will be effect (increase or decrease) of an increase in exogenous income on P*, the equilibrium price?

2) Given the following macroeconomic model.

Y = C + I0 + G0 Equilibrium Income (Y)

C = a + b(Y − T) Consumption with Taxes (T), where: a > 0, 0 < b < 1

T = d + tY Tax function, where: d > 0, 0 < t < 1

Note: I0 = exogenous income, G0 = exogenous govt. spending

Discuss the effect (increase or decrease of an increase in the tax rate, t, on the equilibrium income Y. using the effect of increasing t on the equilibrium solution.
1
Expert's answer
2019-09-06T09:34:56-0400

1.

a)The equilibrium price is the market price where the quantity of goods supplied is equal to the quantity of goods demanded. This is the point at which the demand and supply curves in the market intersect. To determine the equilibrium price, you have to figure out at what price the demand and supply curves intersect.

b)Quantity demanded is a function of price: Qd= f (P). A demand curve can be stated algebraically as Qd= a + bP, where a = intercept, b = slope (which is negative) and P. This is the demand equation, which illustrates the relationship between price and quantity demanded, ceteris paribus .

c) The new equilibrium occurs at a higher quantity and a higher price than the original equilibrium

2. If taxes increase with income, they will also decrease the effect of a change in income on consumption and hence demand. This will dampen the multiplier effect of a change in autonomous expenditure on equilibrium output.


https://www.slideshare.net/nigamvipul17/demand-and-supply-functions-in-economics-66630140